3 Prevailing market price is 4 USD. Firm's total cost and output is given in the following: Output = 0, 10, 20, 30, 40, 50, 60, 65, 70, 75, 80, 85, 90 Total cost = 65, 100, 130, 150, 160, 170, 185, 196, 210, 227, 250, 290, 360 A. Determine the output and dollar amount at which the firm maximizes total profits? B. At what levels of output does the firm break even? (Plot the curves and show graphically with help of total and marginal analysis). C. Determine the output and dollar amount at which the firm maximizes average profits?


C. Determine the output and dollar amount at which the firm maximizes average profits?

How can we calculate part c ?? Average profit=marginal revenue , but both there is a confusion , kindly share your explanation
1
Expert's answer
2020-06-10T18:56:53-0400

Q = 0, 10, 20, 30, 40, 50, 60, 65, 70, 75, 80, 85, 90

TC = 65, 100, 130, 150, 160, 170, 185, 196, 210, 227, 250, 290, 360

MC = - , 3.5, 3.25, 2.83, 2.38, 2.1, 2.0, 2.02, 2.07, 2.16, 2.31, 2.65, 3.28.

A. The output and dollar amount at which the firm maximizes total profits is where MR = MC, so Q = 75 and TP = 4×75 - 227 = 73.

B. The firm break even if P = ATC, it is true if Q = 40.

C. If the price of the goods does not depend on the volume of output of the company, then when the volume of output changes, the price remains unchanged, and the marginal revenue is equal to the revenue from the sale of an additional unit of goods (selling price): MR = AR = P.

MR = MC, so Q = 75 and TP = 4×75 - 227 = 73.